# Enhancing Robustness of Vision-Language Models through Orthogonality Learning and Cross-Regularization

### v1

@@ 3.2 Orthogonal finetuning
where $\bm{W}$ denotes the finetuned weight matrix and $\bm{I}$ is an identity matrix. During the finetuning, we optimize the added $\bm{A}$ while keeping the pretrained weights $\bm{W}_{0}$ frozen. To finetune the model from $\bm{W}_{0}$, we initialize the orthogonal matrix $\bm{A}$ to be identity matrix $\bm{I}$, sharing similar principle with LoRA to set zero initialization of the additive matrices. Moreover, this allows us to gradually inject task-specific knowledge into the finetuned model driven by cross-entropy loss.
Motivated by previous works [51; 42; 28] discussing about differential orthogonalization methods, we focus on taking utilization of Cayley parameterization. The Cayley transform produces a representation of orthogonal matrices without $-1$ eigenvalues using skew-symmetric matrices (i.e., $\bm{C}^{\top}=-\bm{C}$) as follows:
@@
wherein we find this special orthogonal group is able to obtain competitive performances when adapting CLIP for downstream tasks (e.g., few-shot image recognition). Based on the orthogonal finetuning above to adapt the VLM into downsream scenario, we find there exists a potential risky error bounding such that the finetuned model presents inferior generalizability on new/unseen classes, shown in our experimental part. After applying the Neumann series to analyze: $\bm{A}=(\bm{I+C})^{-1}(\bm{I-C})$ can be written as: $\bm{A}\approx\bm{I}+2\bm{C}+\mathcal{O}(\bm{C}^{2})$, We empirically observe that this approximation results in instability of the finetuning [70], which degrades the zero-shot generalization of the pretrained model, showing different phenomena with previous work [63] on finetuning generative models.
Figure 2: Overview of our proposed pipeline, OrthCR. The top shows our finetuning pipeline by applying orthogonal tuning into the Feed-Forward-Network of both image and text encoder ($\mathcal{F}_{v}$ and $\mathcal{F}_{t}$) of CLIP model which is trained with Cross-Regularization strategy. On the left of bottom, orthogonal matrix injection is explained by injecting orthogonal matrix into the pretrained weights with orthogonalization constraint, Cayley parameterization. On the right of bottom, pretrained CLIP is utilized to highlight the most-discriminative image regions and then apply cutout operation to obtain cutout image $X_{cutout}$ which will be input to the finetuned model together with original $X$.
3.3 Cross-regularization
This inspires us to investigate the regularization strategy to carefully constrain the finetuned model not deviating far away from the pretrained one. Therefore, we further design a Cross-Regularization strategy to regularize the finetuned model through pretrained model with a bypass manner since the pretrained weights are frozen. As shown in Fig 2, the text prompts are processed by frozen text encoder $\mathcal{F}_{t}$ to obtain text embedding ${f}_{t}$, while we can also compute new text embedding $f_{t}(:,A_{t})$ which is encoded by orthogonal tuning text encoder after injecting orthogonal matrix to each FFN layer, $\mathcal{F}_{t}+A_{t}$. Here, we want to optimize the additive $A_{t}$ for the text encoder. At the same time, we input original image to the image encoder, and obtain $f_{v}$ encoded by frozen $\mathcal{F}_{v}$ and $f_{v}(:,A_{v})$ from $\mathcal{F}_{v}+A_{v}$, enabling $A_{v}$ tunable only. Further, the pretrained and finetuned logit are computed as follows:
@@ 3.3 Cross-regularization
To further impose regularization from the pretrained anchor point, we first detach the gradients from image encoder to compute new logit $f^{{}^{\prime}}_{logit}$ as:
$f^{{}^{\prime}}_{logit}=sim(f_{t}(:,A_{t})\cdot detach(f_{v}(:,A_{v}))$ (7)
where the text encoder with tunable $A_{t}$ will be optimized to align with the task-specific knowledge (e.g., few-shot image recognition). Then Kullback-Leibler loss $\mathcal{L}_{kl}$ is used to distill informative zero-shot knowledge from the anchor point so as to alleviate deviation far away from the pretrained mainfold wthin a bypass manner, as follows:
$\mathcal{L}_{kl}=\mathcal{D}_{kl}(f^{{}^{\prime}}_{logit},f_{zs\_logit})$ (8)
where $\mathcal{D}_{kl}(f^{{}^{\prime}}_{logit}||f_{zs\_logit})=\sum\limits_{x\in X}(g(f^{{}^{\prime}}_{logit})log{\frac{g(f^{{}^{\prime}}_{logit})}{g(f_{zs\_logit}})})$, $g(\cdot)$ denotes softmax function.
@@ 3.4 Cutout augmentation
As shown in Fig 2, we utilize the pretrained model to infer the similarity map by computing the cosine similarity between image patch tokens and $[CLS]$ text token, which produce a map that each patch responses to $[CLS]$ text token and then reshape them into the same shape of the input image. During the training, we randomly select a cutout region size to zero the top-$K$ image patches, where $K$ ranges from $[l,L]$. To enforce randomness to image encoder so that the model can pay more attention to other less-discriminative image regions, we generate random and different erasing size for each training iteration. Specifically, let $X_{cutout}$ be the cutout image. We input it into the image encoder with $\mathcal{F}_{v}+A_{v}$ and obtain $f_{v\_cutout}(:,A_{v})$. After that, following the aforementioned way, we then calculate the cutout logit $f_{cutout\_logit}$ as:
@@
and newly detached cutout logit $f^{{}^{\prime}}_{cutout\_logit}$ as:
$f^{{}^{\prime}}_{cutout\_logit}=sim(detach(f_{t}(:,A_{t}))\cdot f_{v\_cutout}(:,A_{v}))$ (10)
@@
$\mathcal{L}_{\text{cutout\_ce}}=-\log{p(\hat{y}=y|{X_{cutout}})},\quad\mathcal{L}_{cutout\_kl}=\mathcal{D}_{kl}(f^{{}^{\prime}}_{cutout\_logit},f_{zs\_logit})$ (11)
In this way, we enforce the finetuned model pay more attention to the other less-discriminative image regions that response weak to the text embedding but still contains informative cues to help model learn task-specific knowledge under the data-efficient setting.
@@ 3.5 Training objective
Overall, the training losses of our method consist of two parts, one for the image classification loss including global image classification loss and cutout image classification loss, while the other one includes two corresponding distillation loss. We expect that introducing orthogonal tranformation into CLIP model finetuned for specific downstream tasks is able to retain strong generalizability preservation. Hence, the overall loss $\mathcal{L}_{final}$ can be written as:
@@ 4.1 Experimental settings
Datasets: For evaluation in terms of both base-to-base and base-to-new class generalization, we conduct our method on publicly available 11 image recognition datasets: ImageNet [67] and Caltech101 [19] for generic objects classification, Oxford_Pets [60], StanfordCars [39], Flowers102 [58], Food101 [5] and FGVCAircraft [55] for fine-grained classification, SUN397 [79] for scene recognition, DTD [14] for texture classification, EuroSAT [27] for satellite imagery recognition and UCF101 [71] for action recognition. Following the existing methods [87; 37; 38; 13; 83; 59; 89; 54; 74], we also evaluate our method on cross-dataset transfer and domain generalization. For cross-dataset transfer, we adopt ImageNet as the source and the remaining 10 datasets as target variants, while for domain generalization, we also use ImageNet as source and ImageNetV2 [66], ImageNet-Sketch [75], ImageNet-A [30] and ImageNet-R [29] as targets.
@@
Implementation details: For all the experimental settings, we follow the common strategy of CoOp [88] and CoCoOp [87] for the fair comparison, including the dataset splits, default data augmentation, training schedule, shot of samples, backbones, length of context tokens (i.e., $M$ is 16 in this paper), etc. The $K$ is set to be 3 and averaged for all the experiments, reporting base and novel class accuracy and their harmonic mean (HM), respectively. We apply CLIP-ViT-B/16 as our pretrained backbone model to train for 5 epochs with a batch size of 4, and a learning rate of 1e-5 via SGD optimizer on a single Nvidia-A100-GPU, unless other stated. The hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ are set to be 1.5 and 1.2 by default, left for hyper-parameters sensitivity ablations in Appendix A.
@@
Baseline: To validate the effectiveness of proposed OrthCR, we compare our approach against the following methods, including: (1) zero-shot CLIP [64], which provides the basic baseline model for comparison without any prompt learning or adaptation finetuning; (2) commonly used single-modal prompt tuning methods to demonstrate superiority of our novel finetuning method, such as CoOp [88] which constructs another baseline model for us using tunable context vectors for the input text prompt, CoCoOp [87], PLOT [10] and UNIGRAM [43], and VPT [36]; and multi-modal prompt tuning methods: MaPLe [37] and PromptSRC [38]. Note that the original paper of PLOT [10] adopts a weaker backbone model ResNet-50 [26], here we change it to ViT-B/16 to implement for a fair comparison. Moreover, we also implement VPT which applies prompt tuning for image encoder, IVLP which applies independent prompt tuning for both image encoder and text encoder, and $LoRA_{CLIP}$ that also finetunes the pretrained weights by additive decomposition matrices, all of which establish the basic comparisons.
@@
Table 1: Performance for base-to-base/base-to-new on 11 datasets. We train our model with a subset of the classes (base classes) in a 16-shot setting and evaluate on the test set including base classes and new classes, while HM denotes the harmonic mean of base and novel performance to show the generalization trade-off [78], HM=(2 $\times$ base $\times$ new )/(base + new). The highest results are highlighted in Bold.
@@
Dataset CLIP CoOp CoCoOp MaPLe RPO PLOT PromptSRC UNIGRAM VPT IVLP $LoRA_{CLIP}$ OrthCR Gain
[64] [88] [87] [37] [40] [10] [38] [43] (Base) (Base) (Base) (Ours) $\Delta$
Average on 11 datasets Base 69.34 82.69 80.47 82.28 81.13 77.20 84.26 80.34 80.81 81.83 81.42 84.16 +1.47
New 74.22 63.22 71.69 75.14 75.00 60.38 76.10 75.92 70.36 73.63 70.40 76.55 +13.3
HM 71.70 71.66 75.83 78.55 77.78 67.76 79.97 78.07 74.68 77.10 75.07 80.02 +8.36
ImageNet Base 72.43 76.47 75.98 76.66 76.60 75.97 77.60 76.60 70.93 76.80 74.57 78.10 +1.63
New 68.14 67.88 70.43 70.54 71.57 69.23 70.73 70.69 65.90 70.40 65.50 70.35 +2.47
HM 70.22 71.92 73.10 73.47 74.00 72.44 74.01 73.53 68.32 73.46 69.74 74.02 +2.10
Caltech 101 Base 96.84 98.00 97.96 97.74 97.97 96.53 98.10 98.07 97.86 97.53 98.10 98.17 +0.17
New 94.00 89.81 93.81 94.36 94.37 82.86 94.03 95.11 93.76 93.57 93.25 94.03 +4.22
HM 95.40 93.73 95.84 96.02 96.03 89.17 96.02 96.57 95.77 95.51 95.61 96.06 +2.33
Oxford Pets Base 91.17 93.67 95.20 95.43 94.63 93.45 95.33 94.94 94.81 95.50 94.30 95.60 +1.95
New 97.26 95.29 97.69 97.76 97.50 79.76 97.30 97.94 96.00 97.97 95.30 97.70 +2.41
HM 94.12 94.47 96.43 96.58 96.05 86.06 96.30 96.42 95.40 96.72 94.80 96.64 +2.17
Stanford Cars Base 63.37 78.12 70.49 72.94 73.87 61.41 78.27 73.50 72.46 73.27 73.07 79.40 +1.28
New 74.89 60.40 73.59 74.00 75.53 42.69 74.97 75.38 73.38 74.17 68.53 73.87 +13.4
HM 68.65 68.13 72.01 73.47 74.69 50.37 76.58 74.43 72.92 73.72 70.73 76.54 +8.41
Flowers 102 Base 72.08 97.60 94.87 95.92 94.13 95.62 98.07 95.20 95.39 96.47 95.60 97.60 +0.00
New 77.80 59.67 71.75 72.46 76.67 56.03 76.50 76.21 73.87 72.90 67.60 75.53 +15.8
HM 74.83 74.06 81.71 82.56 84.50 70.56 85.95 84.65 83.26 83.04 79.20 85.16 +11.1
Food101 Base 90.10 88.33 90.70 90.71 90.33 88.45 90.67 90.84 89.88 90.47 87.90 90.50 +0.40
New 91.22 82.26 91.29 92.05 90.83 85.28 91.53 92.12 87.76 91.97 88.93 91.17 +8.91
HM 90.66 85.19 90.99 91.38 90.58 86.84 91.10 91.48 88.81 91.21 88.41 90.83 +5.64
FGVC Aircraft Base 27.19 40.44 33.41 37.44 37.33 29.63 42.73 32.25 33.10 34.20 36.77 41.93 +1.49
New 36.29 22.30 23.71 35.61 34.20 16.17 37.87 38.00 30.49 34.00 31.87 36.87 +14.5
HM 31.09 28.75 27.74 36.50 35.70 20.92 40.15 34.89 31.74 34.10 34.15 39.24 +10.4
SUN397 Base 69.36 80.60 79.74 80.82 80.60 78.56 82.67 80.43 79.66 81.00 79.40 82.47 +1.87
New 75.35 65.89 76.86 78.70 77.80 72.34 78.57 77.91 72.68 78.40 74.47 79.33 +13.4
HM 72.23 72.51 78.27 79.75 79.18 75.32 80.52 79.15 76.01 79.68 76.86 80.87 +8.36
DTD Base 53.24 79.44 77.01 80.36 76.70 69.87 83.37 73.62 79.15 79.50 79.53 82.40 +2.96
New 59.90 41.18 56.00 59.18 62.13 53.63 62.97 62.38 50.76 50.10 52.27 65.33 +24.1
HM 56.37 54.24 64.85 68.16 68.61 60.68 71.75 67.56 61.85 61.47 63.08 72.88 +18.6
EuroSAT Base 56.48 92.19 87.49 94.07 86.63 87.39 92.90 86.26 93.01 91.30 92.67 93.27 +1.08
New 64.05 54.74 60.04 73.23 68.97 67.63 73.90 71.38 54.89 68.53 61.30 79.00 +24.2
HM 60.03 68.69 71.21 82.35 76.79 74.30 82.32 78.12 69.04 78.29 73.89 85.54 +16.8
UCF101 Base 70.53 84.69 82.33 83.00 83.67 72.71 87.10 82.00 82.67 84.13 83.67 86.33 +1.64
New 77.50 56.05 73.45 78.66 75.43 41.51 78.80 78.06 74.54 77.90 75.40 78.87 +22.8
HM 73.85 67.46 77.64 80.77 79.34 52.84 82.74 79.98 78.39 80.90 79.32 82.43 +14.9

### v3

@@ 3.2 Orthogonal fine-tuning
Figure 2: Overview of our proposed pipeline, OrthSR. The top shows our fine-tuning pipeline by applying orthogonal tuning into the Feed-Forward-Network of both image and text encoder ($\mathcal{F}_{v}$ and $\mathcal{F}_{t}$) of CLIP model which is trained with Self-Regularization strategy. On the left of bottom, orthogonal matrix injection is explained by injecting orthogonal matrix into the pretrained weights with orthogonalization constraint (such as Cayley parameterization). On the right of bottom, pretrained CLIP is utilized to highlight the most-discriminative image regions and then apply cutout operation to obtain cutout image $X_{cutout}$ which will be input to the fine-tuned model together with original $X$.
@@
where $\bm{W}$ denotes the fine-tuned weight matrix and $\bm{I}$ is an identity matrix. During the fine-tuning, we optimize the added $\bm{A}$ while keeping the pretrained weights $\bm{W}_{0}$ frozen. To finetune the model from $\bm{W}_{0}$, we initialize the orthogonal matrix $\bm{A}$ to be identity matrix $\bm{I}$, sharing similar principle with LoRA to set zero initialization of the additive matrices. Moreover, this allows us to gradually inject task-specific knowledge into the fine-tuned model driven by cross-entropy loss.
Motivated by previous works Liu et al. (2021); Lezcano-Casado & Martınez-Rubio (2019); Helfrich et al. (2018) discussing about differential orthogonalization methods, we focus on taking utilization of Cayley parameterization. The Cayley transform produces a representation of orthogonal matrices without $-1$ eigenvalues using skew-symmetric matrices (i.e., $\bm{C}^{\top}=-\bm{C}$) as follows:
@@
wherein we find this special orthogonal group is able to obtain competitive performances when adapting CLIP for downstream tasks (e.g., few-shot image recognition). Based on the orthogonal fine-tuning above to adapt the VLM into downsream scenario, we find there exists a potential risky error bounding such that the fine-tuned model presents inferior generalizability on new/unseen classes, shown in our experimental part. After applying the Neumann series to analyze: $\bm{A}=(\bm{I+C})^{-1}(\bm{I-C})$ can be written as: $\bm{A}\approx\bm{I}+2\bm{C}+\mathcal{O}(\bm{C}^{2})$, We empirically observe that this approximation results in instability of the fine-tuning Singla & Feizi (2021), which degrades the zero-shot generalization of the pretrained model, showing different phenomena with previous work Qiu et al. (2023) on fine-tuning generative models.
3.3 Self-Regularization
This inspires us to investigate the regularization strategy to carefully constrain the fine-tuned model not deviating far away from the pretrained one. Therefore, we further design a Self-Regularization strategy to regularize the fine-tuned model through pretrained model with a bypass manner since the pretrained weights are frozen. As shown in Fig 2, the text prompts are processed by frozen text encoder $\mathcal{F}_{t}$ to obtain text embedding ${f}_{t}$, while we can also compute new text embedding $f_{t}(:,A_{t})$ which is encoded by orthogonal tuning text encoder after injecting orthogonal matrix to each FFN layer, $\mathcal{F}_{t}+A_{t}$. Here, we want to optimize the additive $A_{t}$ for the text encoder. At the same time, we input original image to the image encoder, and obtain $f_{v}$ encoded by frozen $\mathcal{F}_{v}$ and $f_{v}(:,A_{v})$ from $\mathcal{F}_{v}+A_{v}$, enabling $A_{v}$ tunable only. Further, the pretrained and fine-tuned logit are computed as follows:
@@ 3.3 Self-Regularization
To further impose regularization from the pretrained anchor point,Then Kullback-Leibler loss $\mathcal{L}_{kl}$ is used to distill informative zero-shot knowledge from the anchor point so as to alleviate deviation far away from the pretrained mainfold wthin a bypass manner, as follows:
$\mathcal{L}_{kd}=\mathcal{D}_{kd}(f_{logit},f_{zs\_logit})$ (7)
where $\mathcal{D}_{kd}(f_{logit}||f_{zs\_logit})=\sum\limits_{x\in X}(g(f_{logit})log{\frac{g(f_{logit})}{g(f_{zs\_logit}})})$, $g(\cdot)$ denotes softmax function.
@@ 3.4 Cutout augmentation
As shown in Fig 2, we utilize the pretrained model to infer the similarity map by computing the cosine similarity between image patch tokens and $[CLS]$ text token, named as attentive CutOut. Then it produces a map that each patch responses to $[CLS]$ text token and then reshape them into the same shape of the input image. During the training, we randomly select a cutout region size to zero the top-$K$ image patches, where $K$ ranges from $[l,L]$. To enforce randomness to image encoder so that the model can pay more attention to other less-discriminative image regions, we generate random and different erasing size for each training iteration. Specifically, let $X_{cutout}$ be the cutout image. We input it into the image encoder with $\mathcal{F}_{v}+A_{v}$ and obtain $f_{v\_cutout}(:,A_{v})$. After that, following the aforementioned way, we then calculate the cutout logit $f_{cutout\_logit}$ as:
@@
$\mathcal{L}_{\text{cutout\_ce}}=-\log{p(\hat{y}=y|{X_{cutout}})},\quad\mathcal{L}_{cutout\_kd}=\mathcal{D}_{kd}(f_{cutout\_logit},f_{zs\_logit})$ (9)
In this way, we enforce the fine-tuned model pay more attention to other less-discriminative image regions that response weak to the text embedding but still contains informative cues to help model learn task-specific knowledge under the data-efficient setting, which serves as diversifying samples.
@@ 3.5 Training objective
Overall, the training losses of our method consist of two parts, one for the image classification loss including global image classification loss and cutout image classification loss, while the other one includes two corresponding distillation loss. We expect that introducing orthogonal tranformation into CLIP model fine-tuned for specific downstream tasks is able to retain strong generalizability preservation. Hence, the overall loss $\mathcal{L}_{final}$ can be written as:
$\mathcal{L}_{\text{final}}=\lambda_{1}(\mathcal{L}_{ce}+\mathcal{L}_{cutout\_ce})+\lambda_{2}(\mathcal{L}_{kd}+\mathcal{L}_{cutout\_kd})$ (10)
@@
3.6 Theoretical Analysis
In this section, we provide theoretical analysis for the generalization error bound of  OrthSR.
We define the following optimization objectives according to Eq. 10:
$\min_{\Theta\in\mathbb{R}}\underset{\mathcal{L}_{CE}}{\underbrace{\frac{1}{N}{\textstyle\sum_{i=1}^{N}}\mathcal{L}\left(\hat{s}_{i}^{S}\left(\Theta\right),y_{i}^{gt}\right)}}+\lambda\underset{\mathcal{L}_{KD}}{\underbrace{\mathcal{L}\left(\hat{s}^{S}\left(\Theta\right),\hat{s}^{T}\right)}},\vspace{0pt}$ (11)
where $\Theta$ represents learnable orthogonal matrices $\{A_{v},A_{t}\}$ of the proposed method, and we use $S$ and $T$ here to denote the fine-tuned model and pre-trained  anchor model. Now we further analyze the effectiveness of  OrthSR by computing the generalization error bound. This bound computes the bias between the generalization error $\varepsilon\left(\Theta\right):=\bm{\mathbb{E}}_{\left(\hat{s}^{S},y^{gt}\right)\sim\mathcal{D}}\mathcal{L}\left(\hat{s}^{S}\left(\Theta\right),y^{gt}\right)$ and empirical error $\bar{\varepsilon}_{\chi}\left(\Theta\right):=\frac{1}{N}{\textstyle\sum_{i=1}^{N}}\mathcal{L}\left(\hat{s}_{i}^{S}\left(\Theta\right),y_{i}^{gt}\right)$, where $D$ is the real data distribution and $\bm{\mathbb{E}}\left(\cdot\right)$ denotes the expectation function.
Theorem 1.
Assume that $\Theta^{*}$ is the solution to Eq. equation 11. Then we have that for any $10<\epsilon<$ 1 with probability $1-\epsilon$,
$\epsilon(\Theta^{*})-\bar{\epsilon}_{\chi}(\Theta^{*})\leq X^{*}\sqrt{\frac{2\ln(1/\delta)}{N}}+\frac{C^{\prime\prime}}{\lambda^{2\alpha}\sqrt{N}}.$
where $X^{*}=\mathrm{max}_{r\in\mathbb{N}_{N}}\left|\mathcal{L}\left(\hat{s}_{r}^{S}\left(\Theta\right),y_{r}^{gt}\right)\right|$ and $\alpha>0$.
The first term of the upper bound converges with the increasing of the number of training data $N$, that can be achieved by our proposed attentive CutOut data augmentation instead of using extra data. We can also find that the second term converges to 0 with the increasing of $\lambda$, which means the our self-regularization $\mathcal{L}_{KD}$ within a bypass manner effectively improves the generalization ability of our method.
@@ 4.1 Experimental settings
Datasets: For evaluation in terms of both base-to-base and base-to-new class generalization, we conduct our method on publicly available 11 image recognition datasets: ImageNet Russakovsky et al. (2015) and Caltech101 Fei-Fei (2004) for generic objects classification, Oxford_Pets Parkhi et al. (2012), StanfordCars Krause et al. (2013), Flowers102 Nilsback & Zisserman (2008), Food101 Bossard et al. (2014) and FGVCAircraft Maji et al. (2013) for fine-grained classification, SUN397 Xiao et al. (2010) for scene recognition, DTD Cimpoi et al. (2014) for texture classification, EuroSAT Helber et al. (2019) for satellite imagery recognition and UCF101 Soomro et al. (2012) for action recognition. Following the existing methods Zhou et al. (2022a); Khattak et al. (2023a; b); Cho et al. (2023); Zang et al. (2022b); Park et al. (2024); Zhu et al. (2023); Lu et al. (2022); Wang et al. (2023), we also evaluate our method on cross-dataset transfer and domain generalization. For cross-dataset transfer, we adopt ImageNet as the source and the remaining 10 datasets as target variants, while for domain generalization, we also use ImageNet as source and ImageNetV2 Recht et al. (2019), ImageNet-Sketch Wang et al. (2019), ImageNet-A Hendrycks et al. (2021b) and ImageNet-R Hendrycks et al. (2021a) as targets.
@@
Implementation details: For all the experimental settings, we follow the common strategy of CoOp Zhou et al. (2022b) and CoCoOp Zhou et al. (2022a) for the fair comparison, including the dataset splits, default data augmentation, training schedule, shot of samples, backbones, length of context tokens (i.e., $M$ is 16 in this paper), etc. The $K$ is set to be 3 and averaged for all the experiments, reporting base and novel class accuracy and their harmonic mean (HM), respectively. We apply CLIP-ViT-B/16 as our pretrained backbone model to train for 5 epochs with a batch size of 4, and a learning rate of 1e-5 via SGD optimizer on a single Nvidia-A100-GPU, unless other stated. The hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ are set to be 1.5 and 1.2 by default, left for hyper-parameters sensitivity ablations in Appendix A.
@@
Baseline: To validate the effectiveness of proposed OrthSR, we compare our approach against the following methods, including: (1) zero-shot CLIP Radford et al. (2021), which provides the basic baseline model for comparison without any prompt learning or adaptation finetuning; (2) commonly used single-modal prompt tuning methods to demonstrate superiority of our novel finetuning method, such as CoOp Zhou et al. (2022b) which constructs another baseline model for us using tunable context vectors for the input text prompt, CoCoOp Zhou et al. (2022a), PLOT Chen et al. (2023) and UNIGRAM Li et al. (2023a), and VPT Jia et al. (2022); and multi-modal prompt tuning methods: MaPLe Khattak et al. (2023a) and PromptSRC Khattak et al. (2023b). Note that the original paper of PLOT Chen et al. (2023) adopts a weaker backbone model ResNet-50 He et al. (2016), here we change it to ViT-B/16 to implement for a fair comparison. Moreover, we also implement VPT which applies prompt tuning for image encoder, IVLP which applies independent prompt tuning for both image encoder and text encoder, all of which establish the basic comparisons.
@@
Table 1: Performance for base-to-base/base-to-new on 11 datasets. We train our model with a subset of the classes (base classes) in a 16-shot setting and evaluate on the test set including base classes and new classes, while HM denotes the harmonic mean of base and novel performance to show the generalization trade-off Xian et al. (2017), HM=(2 $\times$ base $\times$ new )/(base + new). The highest results are highlighted in Bold.
@@
Dataset CLIP CoOp CoCoOp MaPLe RPO PLOT PromptSRC UNIGRAM VPT IVLP OrthSR Gain
Radford et al. (2021) Zhou et al. (2022b) Zhou et al. (2022a) Khattak et al. (2023a) Lee et al. (2023) Chen et al. (2023) Khattak et al. (2023b) Li et al. (2023a) (Base) (Base) (Ours) $\Delta$
Average on 11 datasets Base 69.34 82.69 80.47 82.28 81.13 77.20 84.26 80.34 80.81 81.83 84.16 +1.47
New 74.22 63.22 71.69 75.14 75.00 60.38 76.10 75.92 70.36 73.63 76.55 +13.3
HM 71.70 71.66 75.83 78.55 77.78 67.76 79.97 78.07 74.68 77.10 80.02 +8.36
ImageNet Base 72.43 76.47 75.98 76.66 76.60 75.97 77.60 76.60 70.93 76.80 78.10 +1.63
New 68.14 67.88 70.43 70.54 71.57 69.23 70.73 70.69 65.90 70.40 70.35 +2.47
HM 70.22 71.92 73.10 73.47 74.00 72.44 74.01 73.53 68.32 73.46 74.02 +2.10
Caltech 101 Base 96.84 98.00 97.96 97.74 97.97 96.53 98.10 98.07 97.86 97.53 98.17 +0.17
New 94.00 89.81 93.81 94.36 94.37 82.86 94.03 95.11 93.76 93.57 94.03 +4.22
HM 95.40 93.73 95.84 96.02 96.03 89.17 96.02 96.57 95.77 95.51 96.06 +2.33
Oxford Pets Base 91.17 93.67 95.20 95.43 94.63 93.45 95.33 94.94 94.81 95.50 95.60 +1.95
New 97.26 95.29 97.69 97.76 97.50 79.76 97.30 97.94 96.00 97.97 97.70 +2.41
HM 94.12 94.47 96.43 96.58 96.05 86.06 96.30 96.42 95.40 96.72 96.64 +2.17
Stanford Cars Base 63.37 78.12 70.49 72.94 73.87 61.41 78.27 73.50 72.46 73.27 79.40 +1.28
New 74.89 60.40 73.59 74.00 75.53 42.69 74.97 75.38 73.38 74.17 73.87 +13.4
HM 68.65 68.13 72.01 73.47 74.69 50.37 76.58 74.43 72.92 73.72 76.54 +8.41
Flowers 102 Base 72.08 97.60 94.87 95.92 94.13 95.62 98.07 95.20 95.39 96.47 97.60 +0.00
New 77.80 59.67 71.75 72.46 76.67 56.03 76.50 76.21 73.87 72.90 75.53 +15.8
HM 74.83 74.06 81.71 82.56 84.50 70.56 85.95 84.65 83.26 83.04 85.16 +11.1
Food101 Base 90.10 88.33 90.70 90.71 90.33 88.45 90.67 90.84 89.88 90.47 90.50 +0.40
New 91.22 82.26 91.29 92.05 90.83 85.28 91.53 92.12 87.76 91.97 91.17 +8.91
HM 90.66 85.19 90.99 91.38 90.58 86.84 91.10 91.48 88.81 91.21 90.83 +5.64
FGVC Aircraft Base 27.19 40.44 33.41 37.44 37.33 29.63 42.73 32.25 33.10 34.20 41.93 +1.49
New 36.29 22.30 23.71 35.61 34.20 16.17 37.87 38.00 30.49 34.00 36.87 +14.5
HM 31.09 28.75 27.74 36.50 35.70 20.92 40.15 34.89 31.74 34.10 39.24 +10.4
SUN397 Base 69.36 80.60 79.74 80.82 80.60 78.56 82.67 80.43 79.66 81.00 82.47 +1.87
New 75.35 65.89 76.86 78.70 77.80 72.34 78.57 77.91 72.68 78.40 79.33 +13.4
HM 72.23 72.51 78.27 79.75 79.18 75.32 80.52 79.15 76.01 79.68 80.87 +8.36
DTD Base 53.24 79.44 77.01 80.36 76.70 69.87 83.37 73.62 79.15 79.50 82.40 +2.96
New 59.90 41.18 56.00 59.18 62.13 53.63 62.97 62.38 50.76 50.10 65.33 +24.1
HM 56.37 54.24 64.85 68.16 68.61 60.68 71.75 67.56 61.85 61.47 72.88 +18.6
EuroSAT Base 56.48 92.19 87.49 94.07 86.63 87.39 92.90 86.26 93.01 91.30 93.27 +1.08
New 64.05 54.74 60.04 73.23 68.97 67.63 73.90 71.38 54.89 68.53 79.00 +24.2
HM 60.03 68.69 71.21 82.35 76.79 74.30 82.32 78.12 69.04 78.29 85.54 +16.8
UCF101 Base 70.53 84.69 82.33 83.00 83.67 72.71 87.10 82.00 82.67 84.13 86.33 +1.64
New 77.50 56.05 73.45 78.66 75.43 41.51 78.80 78.06 74.54 77.90 78.87 +22.8
HM 73.85 67.46 77.64 80.77 79.34 52.84 82.74 79.98 78.39 80.90 82.43 +14.9